5. Solve the initial value problem y + 4y = 2 + 3e, y(0) = 0, y(0) = 2.
Answer #1

y + 4y = + + 3et y (o)=0.3 (0)=2 Applying Laplace on both sides Ily] + 4L[y] = { [ta+ Bet) sycol-y(o) S²YG) -2= 2/3 8² x (
Answer #2

y"+4y´+4y=3e-t  y(0)=2;y´(0)=1